A 0.2025 g sample consisting of only BaCl2

and KCl required 20.25 mL of 0.1200M AgNO3 solution for the quantitative precipitation of chloride.Calculate the %Ba and %K in the sample


1
Expert's answer
2021-02-09T03:46:54-0500

Let y g of Kcl and ( 0.2 - y ) g of BaCl2 .


y / 74.55 + ( 0.2 - y ) × 2 / 208.23


= 20.25 × 0.12 / 1000.


Then we can find y .

Then % of Kcl = y × 100 / 0.2025

Then % of Bacl2 = ( 0.2025 -y ) × 100 / 0.2025